1.   Calculate the number of grams of helium in a 1.00 L balloon at 27.0∘∘C and under 1.00 atmosphere of pressure.



1
Expert's answer
2021-11-06T01:31:58-0400

PV= nRT

1×1000=n×8.314×300

1000=2494.2n

n= 0.401

Molar mass of helium=4

= 0.401×4=1.604g